### v1

@@ Abstract
keywords:
cavitation |||| bubble nucleation |||| contact angle |||| molecular dynamics simulation
@@
When water is brought into a metastable state by either heating or pressure reduction, it will eventually reach equilibrium by cavitation, i.e., by the nucleation of vapor bubbles. Cavitation and the collapse of cavitation bubbles are central to applications such as inkjet printing, ultrasonic cleaning, noninvasive destruction of kidney stones by acoustic shockwaves, sonochemistry and sonoluminescence (1, 2). In contrast, the collapse of hydrodynamically produced cavitation bubbles has adverse effects on hydraulic systems, valves and propeller blades, where it causes wear and erosion (3, 4, 5). In biology, where temperature is fixed by the environment, metastable water is primarily produced by pressure reduction. The collapse of these cavitation bubbles provides catapult-like mechanisms in ferns (6) and allows snapping shrimp to stun their prey (7). Conversely, cavitation limits the negative pressure produced by octopus suckers (8) and must be avoided in the ascending sap of tall trees, where significantly negative pressures are encountered, in order to prevent embolies (9). Importantly, metastable water at negative pressures can persist over astonishingly long times and cavitation occurs at experimentally relevant time scales only when the pressure sinks below a threshold cavitation pressure $p_{\mathrm{cav}}$, which can be as low as −100 MPa in certain systems (2). The large variation of $p_{\mathrm{cav}}$ among different systems and its dependence on surface properties suggests that cavitation in bulk water (i.e. homogeneous nucleation) competes with cavitation on surfaces (i.e. heterogeneous nucleation) and with cavitation at surface defects. Understanding the kinetic competition between these three different cavitation pathways is important for all the above-mentioned applications and is the central theme of this paper.
@@
Cavitation is a typical nucleation phenomenon; according to Classical Nucleation Theory (CNT), its kinetics is determined by the thermally activated stochastic crossing of a free energy barrier characteristic of the critical nucleus size  (10). CNT predicts that clean water at room temperature can sustain negative pressures considerably below −100 MPa before bulk cavitation sets in (11, 12, 2, 13, 14). This very low cavitation pressure predicted for clean bulk water is largely due to the high water surface tension. Experimentally, cavitation pressures below −100 MPa have only been reached with ultra-clean water in microscopic quartz cavities (15, 16, 17). In typical experiments with hydrophilic surfaces and using rigorous purification and degassing methods, cavitation pressures of the order of $p_{\mathrm{cav}}\approx-30$ MPa (2, 13, 18) are reached, significantly above the CNT results for clean bulk water. However, with untreated ordinary water in standard containers, the cavitation pressure is positive and typically only slightly below the saturated water vapor pressure of about 2 kPa. The consensus in the field is that surface cavitation at container walls, microscopic impurities (19, 20, 21, 22, 23, 24, 25), surface imperfections, or crevices hosting pre-existing vapor or gas bubbles (26, 27, 28) contribute to deviations in cavitation pressures from those predicted by CNT for clean bulk water (25, 29). However, a unified theory that accounts for the simultaneous cavitation in bulk water, at surfaces and at surface defects is missing and is formulated in this paper.
Our kinetic cavitation model predicts the cavitation pressure $p_{\mathrm{cav}}$ as a function of observation time and container size made of surfaces with varying water contact angles and defects with different size and wetting properties. We find that only for very hydrophilic and defect-free surfaces cavitation occurs in bulk and at a cavitation pressure $p_{\mathrm{cav}}$ given by bulk CNT theory. For defect-free surfaces and above a critical surface contact angle $\theta$ between 50° and 60° (depending on systems size and observation time), cavitation instead predominantly starts at the surface. Finally, for surfaces that contain defects that are more hydrophobic than the surface, cavitation proceeds from one of the defects with a rate that depends only weakly on the number or surface concentration of defects. For the quantitative comparison of bulk, surface and defect cavitation kinetics we determine their attempt frequencies from all-atom force-field molecular dynamics (MD) simulations, which in the past have proven useful for the study of cavitation processes (30, 31, 32, 33, 25).
Results and Discussion
@@
0.1 Kinetic model for cavitation in water confined by surfaces with defects
@@
In the typical cavitation scenario, water is confined by solid surfaces with a given contact angle that contain a finite number of surface defects. Upon application of a negative pressure, cavitation can proceed in bulk, at the defect-free surface parts or at the defects. Since we cannot presuppose the prevalence of one cavitation pathway over the others, we write down the total cavitation rate as the sum of the bulk (3D), surface (2D) and defect (def) rates as
@@
where $N_{\mathrm{def}}$ is the number of surface defects. The model assumes that the different cavitation pathways do not interfere with each other, which we confirm by our MD simulations. It will turn out that depending on the surface and defect properties, quantified by their respective contact angles, one cavitation pathway will always dominate strongly over the other pathways, such that we can construct effective kinetic phase diagrams that feature regions where bulk, surface or defect cavitation dominates.
0.2 Cavitation in bulk
CNT is a general framework for the description of cavitation (10). For cavitation in bulk, one considers a spherical bubble in a liquid with a free energy given by (2, 14, 34)
@@ 0.2 Cavitation in bulk
where $\gamma$ is the liquid–vapor surface tension, $r$ the bubble radius and $p<0$ the negative pressure in the liquid. The first term represents the bubble surface free energy and the second term the work associated with the volume expansion. We have neglected the saturated vapor pressure, which is irrelevant compared to the large magnitude of cavitation pressures considered here. For simplicity, we have also neglected curvature effects on the surface tension, which only introduce minor corrections (30, 35). The interplay between the two terms in Eq. 2 creates a free energy barrier for $p<0$ of
@@
$k_{\mathrm{3D}}=\kappa_{\mathrm{3D}}Ve^{-\beta G^{*}_{\mathrm{3D}}}\,.$ (4)
The cavitation time follows from the rate as $\tau=k_{\mathrm{3D}}^{-1}$ (36) and is the mean time one has to wait for cavitation to occur. The exponential factor in Eq. 4 reflects the Arrhenius law (36), where $\beta=1/k_{\mathrm{B}}T$ is the inverse thermal energy. The pre-exponential factor splits into the volume $V$ and the attempt frequency density $\kappa_{\mathrm{3D}}$, which is an intensive property of the liquid that characterizes the barrier-less cavitation rate. It thus constitutes the cavitation speed limit, i.e., the maximum rate at which cavitation can occur (37). The attempt frequency density, $\kappa_{\mathrm{3D}}$, accounts for viscosity and hydrodynamic memory effects in the surrounding liquid, as well as for corrections to CNT, such as the curvature dependence of the interfacial tension (38). Given the exponential dependence of the cavitation time $\tau$ on $p$, it is preferred to consider the inverse relation. For a cubic volume $V=L^{3}$, the cavitation pressure follows from Eqs. 3 and 4 as  (2, 39)
@@
It is seen that $p_{\mathrm{3Dcav}}$ depends only logarithmically and thus rather weakly on system size $L$ and cavitation time $\tau$. Various methods to determine $\kappa_{\mathrm{3D}}$ exist (40, 41). Here, we use the value of $\kappa_{\mathrm{3D}}$ from our previous MD simulations (35), which will allow us to compare with surface-cavitation and defect-cavitation rates obtained by the same method on the same footing. In our previous simulation study on bulk cavitation, we determined $\kappa_{\mathrm{3D}}=5\times 10^{11}\,\mathrm{ns^{-1}nm^{-3}}$ for the SPC/E water model (35), yielding $p_{\mathrm{3Dcav}}$ between --80 and --100 MPa for $L$ between 1 µm and 1 m and $\tau=1$ s according to Eq. 5. This value is slightly lower than previous predictions, primarily because the surface tension of SPC/E water ($\gamma_{\textrm{SPC/E}}\approx 55$ mN/m for a Lennard-Jones cutoff 0.9 nm (42)) is lower than the experimental value $\gamma_{\textrm{exp}}=72$ mN/m. Applying a correction factor of $(\gamma_{\textrm{exp}}/\gamma_{\textrm{SPC/E}})^{3/2}\approx 1.5$ on $p_{\mathrm{3Dcav}}$, as suggested by Eq. 5, yields good agreement with previous predictions (11, 12, 2, 13, 14). We here stick to the SCP/E water model because it reproduces most dynamic water properties quite well.
0.3 Cavitation at a smooth defect-free surface
@@
which is reduced compared to the bulk result $G_{\mathrm{3D}}^{*}$ in Eq. 3 by the geometric factor (43, 10, 2, 21)
@@
Figure 2: (A) Time-dependent pressure ($p$, top) and simulation box size ($L_{z}$, bottom) during a pressure ramp simulation with a pressure rate of $\dot{p}=-5\,\mathrm{MPa/ns}$ for a system with a surface contact angle of $\theta=97^{\circ}$. The presented trajectory corresponds to the final 200 ps prior to cavitation, identified as a 50 % increase of $L_{z}$ relative to its initial value $L_{z}^{0}=8\,\mathrm{nm}$. We find a cavitation pressure of $p_{\mathrm{cav}}^{*}=-93.9\,\mathrm{MPa}$. Vertical lines mark individual simulation segments at constant pressure. The inset shows consecutive snapshots of the cavitation event. (B) Cavitation configurations with $L_{z}/L_{z}^{0}=3/2$ for different contact angles. (C) Simulation results for $(p_{\mathrm{cav}}^{*})^{-2}$ versus $\dot{p}$ for different contact angles. The lines are fits of Eq. 16 to the data, where $\kappa_{\textrm{2D}}$ is the only fitting parameter. (D) Results for $\kappa_{\textrm{2D}}$ (left scale) and the ratio $\xi=\kappa_{\mathrm{2D}}/\kappa_{\mathrm{3D}}$ (right scale) as a function of contact angle $\theta$. The solid black line is a fit of Eq. S1 in the SI Appendix, section S1.
@@
To determine $\kappa_{\textrm{2D}}$ from simulations, we use a self-assembled monolayer of alkyl molecules terminated by headgroups with adjustable electric dipole moments (44, 42). The dipole moment controls the wetting characteristics of the surfaces. We examine six different dipole moments, resulting in contact angles of $45^{\circ}$, $60^{\circ}$, $76^{\circ}$, $97^{\circ}$, $120^{\circ}$ and $135^{\circ}$, on surfaces measuring 9 nm $\times$ 10.4 nm in contact with water. Obtaining the cavitation time from simulations at constant negative pressure is impractical due to the excessive waiting time for cavitation to occur. Instead, we apply negative pressures that decrease linearly over time as $p(t)=\dot{p}t$ with different fixed pressure ramp rates $\dot{p}<0$, as used before for bulk cavitation  (35) and described in the methods section.
@@
Fig. 2 summarizes our simulation results. Panel A illustrates a cavitation event on a hydrophobic surface with contact angle $\theta=97^{\circ}$, where a vapor bubble forms at the surface (see inset snapshots) due to the progressively decreasing pressure over time (upper graph). The simulation trajectory consists of a sequence of 20-ps-long simulations (delineated by gray vertical lines), each at a constant applied pressure that decreases in discrete steps of 0.1 MPa, which corresponds to a mean pressure rate of $\dot{p}=-5\,\mathrm{MPa/ns}$. The formation of the bubble is monitored by the box size normal to the surface $L_{z}$ (bottom graph). We define cavitation to take place when the box size $L_{z}$ increases by 50 % compared to its initial value, which in this specific example yields a cavitation pressure $p_{\mathrm{2Dcav}}^{*}=-93.9\,\mathrm{MPa}$. It is important to note that the cavitation pressure in the pressure-ramp ensemble $p_{\mathrm{2Dcav}}^{*}$ is different from the cavitation pressure $p_{\mathrm{2Dcav}}$ in the constant-pressure ensemble given in Eq. 10, which will be determined later.

### v2

@@ Abstract
preprint: AIP/123-QED
I Introduction
@@
When water is brought into a metastable state by either heating or pressure reduction, it will eventually reach equilibrium by cavitation, i.e., by the nucleation of vapor bubbles. Cavitation and the collapse of cavitation bubbles are central to applications such as inkjet printing, ultrasonic cleaning, noninvasive destruction of kidney stones by acoustic shockwaves, sonochemistry, sonoluminescence Coleman and Saunders (1989); Caupin and Herbert (2006) and for the investigation of geothermal processesFall, Rimstidt, and Bodnar (2009). In contrast, the collapse of hydrodynamically produced cavitation bubbles has adverse effects on hydraulic systems, valves, and propeller blades, causing wear and erosion Dular et al. (2004); Reuter, Deiter, and Ohl (2022); Adhikari, Goliaei, and Berkowitz (2015), while the creation of cavitation bubbles may impact the performance of hydrofoils Custodio, Henoch, and Johari (2018). In biology, where temperature is fixed by the environment, metastable water is primarily produced by pressure reduction. The collapse of these cavitation bubbles provides catapult-like mechanisms in ferns Noblin et al. (2012) and allows snapping shrimp to stun their prey Verslius et al. (2000). Conversely, cavitation limits the negative pressure produced by octopus suckers Smith (1991) and must be avoided in the ascending sap of tall trees, where significantly negative pressures are encountered, in Importantly, metastable water at negative pressures can persist over astonishingly long times and cavitation occurs at experimentally relevant time scales only when the pressure sinks below a threshold cavitation pressure $p_{\mathrm{cav}}$. The large variation of $p_{\mathrm{cav}}$ among different systems and its dependence on surface properties suggests that cavitation in bulk water (i.e., homogeneous nucleation) competes with cavitation on surfaces and with cavitation at surface defects (i.e., heterogeneous nucleation). Understanding the kinetic competition between these three different cavitation pathways is important for all the above-mentioned applications and is the central theme of this paper.
@@
Cavitation is a typical nucleation phenomenon; according to Classical Nucleation Theory (CNT), its kinetics is determined by the thermally activated stochastic crossing of a free energy barrier characteristic of the critical nucleus size Debenedetti (1996). CNT predicts that clean water at room temperature can sustain negative pressures considerably below −100 MPa before bulk cavitation sets in Fisher (1948); Caupin (2005); Caupin and Herbert (2006); Caupin and Stroock (2013); Azouzi et al. (2013a). This very low cavitation pressure predicted for clean bulk water is largely due to the high water surface tension. Experimentally, cavitation pressures below −100 MPa have only been reached with ultra-clean water in microscopic quartz cavities Zheng et al. (1991); Alvarenga, Grimsditch, and Bodnar (1993); Azouzi et al. (2013b). In typical experiments with hydrophilic surfaces and using rigorous purification and degassing methods, cavitation pressures of the order of $p_{\mathrm{cav}}\approx-30$ MPa Caupin and Herbert (2006); Caupin and Stroock (2013); Caupin (2015) are reached, significantly above the CNT results for clean bulk water. However, with untreated ordinary water in standard containers, the cavitation pressure is positive and typically only slightly below the saturated water vapor pressure of about 2 kPa. The consensus in the field is that surface cavitation at container walls, microscopic impurities Jones, Evans, and Galvin (1999); Marschall et al. (2003); Sear (2007); Mørch (2007); Tsuda, Takagi, and Matsumoto (2008); Groß and Pelz (2017); Gao, Wu, and Wang (2021), surface imperfections, or crevices hosting pre-existing vapor or gas bubbles Harvey et al. (1944); Atchley and Prosperetti (1989); Borkent et al. (2009); Pfeiffer et al. (2022) act as cavitation nuclei. These nuclei promote cavitation and contribute to deviations in cavitation pressures from those predicted by CNT for clean bulk water.
These experimental complications make theoretical and simulation approaches essential. Molecular dynamics (MD) simulations with all-atom force fields have proven to be powerful tools for investigating cavitation processes Menzl et al. (2016); P Lamas et al. (2023); Min and Berkowitz (2019); Xie, Xu, and Zhong (2022). Recent MD studies demonstrated that while dissolved gas has a negligible effect on cavitation thresholds Zhou et al. (2019); Kanduč et al. (2020), gas nanobubbles can significantly promote cavitation Dockar, Borg, and Reese (2019); Gao, Wu, and Wang (2021). Indeed, our recent MD study demonstrated that nonsoluble impurities in the form of nanodroplets suspended in water considerably increase cavitation pressures Šako, Netz, and Kanduč (2024). Another simulation study has shown that solid particles in water can disrupt the hydrogen-bond network and facilitate cavitation in small simulation volumes Li, Gu, and Chen (2019). Despite these advances, the influence of specific surface properties on cavitation in macroscopic systems remains poorly understood, and a unified framework encompassing bulk, surface, and defect-induced cavitation is still missing.
In this work, we bridge this gap by formulating a kinetic cavitation model that combines MD simulations with the CNT framework. With attempt rates obtained from simulations, our kinetic model predicts the cavitation pressure, $p_{\mathrm{cav}}$, of water confined in a cubic container, with walls characterized by the water contact angle. The model also incorporates various surface defect patches with distinct wetting properties. We find that for highly hydrophilic, defect-free surfaces, cavitation occurs in bulk water at pressures consistent with CNT predictions. As the surface contact angle increases beyond a critical threshold of 50° – 60°, cavitation shifts to the surface. Finally, for surfaces with hydrophobic defects our comprehensive model provides a unified framework to understand and predict cavitation behavior in diverse practically relevant scenarios.
II Materials and Methods
Simulation model
Figure 1: Simulation setup consisting of 800 surface molecules (hydroxylated alkanes) assembled into two self-assembled monolayers and solvated by 16,353 water molecules. Each surface molecule consists of an alkyl chain terminated by a modified hydroxyl group with its partial charges scaled by a factor $\alpha$. Simulation box (black frame) is replicated in all three directions via periodic boundary conditions. The box length for the initial pressure of $-30\,\mathrm{MPa}$ in the normal direction is $L_{z}^{0}=8\,\mathrm{nm}$.
The simulation setup follows our previous approach Kanduč, Schneck, and Netz (2014); Kanduč and Netz (2017) and is shown in Fig. 1. We model a planar surface composed of two self-assembled monolayers, each consisting of ten-carbon-atom alkyl chains terminated by polar a hydroxyl ($\mathrm{OH}$) head group. These hydroxylated alkanes are arranged on a hexagonal lattice with an areal density of $4.3\,\mathrm{nm^{-2}}$. To stabilize the structure, we apply harmonic restraints to the second carbon atom from the head group with strengths of $k_{x}=k_{y}=500\,\mathrm{(kJ/mol)\,nm^{-2}}$ in the lateral directions, and to the tenth (terminal) carbon atom with $k_{x}=k_{y}=10\,\mathrm{(kJ/mol)\,nm^{-2}}$ laterally and $k_{z}=100\,\mathrm{(kJ/mol)\,nm^{-2}}$ in the normal direction. The surface polarity and thereby the surface contact angle is controlled by scaling the partial charges of the head groups with a factor $\alpha$, ranging from 0 to 1. For $\alpha=0$, the surface is non-polar with head groups resembling methylated termini, resulting in a contact angle of $\theta=135^{\circ}$. For $\alpha=1$, the surface is fully polar, mimicking hydroxyl head groups. In the Supplementary Material, section S1, we present a phenomenological fit function that maps the surface polarity $\alpha$ to the contact angle $\cos\theta$. The periodic simulation box, with initial dimensions $9\,\mathrm{nm}\times 10.4\,\mathrm{nm}\times 8\,\mathrm{nm}$ at a pressure of $-30\,\mathrm{MPa}$, contains 16,353 water molecules and 800 hydroxylated alkanes.
We use united-atom parameters from the GROMOS force field for the alkanes Oostenbrink et al. (2004); Kanduč and Netz (2017) and the rigid SPC/E water model Berendsen, Grigera, and Straatsma (1987). Classical non-polarizable molecular dynamics simulations are performed with the GROMACS package (Abraham et al., 2015) using a time step of $2\,\mathrm{fs}$. A plain cutoff of $0.9\,\mathrm{nm}$ is used for short-range Lennard-Jones interactions. Electrostatics is treated with the Smooth Particle–Mesh Ewald method, using a $0.9\,\mathrm{nm}$ real-space cutoff. A temperature of 300 K is controlled using the velocity rescale thermostat with a stochastic factor Bussi, Donadio, and Parrinello (2007) and a time constant of $0.5\leavevmode\nobreak\ \mathrm{ps}$. Pressure is controlled by the Berendsen barostat with a time constant of $1\,\mathrm{ps}$, suitable even for negative pressures because of its efficiency and stability in box scaling under large pressure differences. We previously verified Kanduč et al. (2020) that equivalent results are obtained with the Parrinello–Rahman barostat, which more accurately reproduces volume fluctuations. Reference coordinates for surface-atom restraints are scaled with the coupling matrix of the pressure coupling.
Kinetic theory for pressure ramps
To obtain the cavitation attempt frequencies, we use the pressure-ramp simulation protocol Kanduč et al. (2020), where the negative pressure decreases linearly over time as $p(t)=\dot{p}t$ at a constant rate $\dot{p}<0$. To simulate pressure ramps, we conduct a sequence of simulations at constant pressures, starting from an initial pressure of $-30\,\mathrm{MPa}$. After each simulation, the final coordinates are used to start the next simulation, with the pressure reduced by $0.1\,\mathrm{MPa}$. The duration of each constant-pressure simulation depends on the pressure rate $\dot{p}$. We identify cavitation when the simulation box length $L_{z}$ increases by more than 50 % of its initial value $L_{z}^{0}$. The corresponding “dynamic” cavitation pressure $p^{*}_{\mathrm{cav}}$ follows from the solution of the linear cavitation rate equation as
$p^{*}_{\mathrm{cav}}=\dot{p}\int_{0}^{\infty}e^{-k_{0}I(t)}\mathrm{d}t\,,$ (1)
where
[…]
Here, $\beta=1/k_{\mathrm{B}}T$ is the inverse thermal energy, and $G^{*}(t)$ is the time-dependent free energy barrier in the pressure-ramp protocol. According to CNT, this barrier decays with time as $G^{*}(t)\sim t^{-2}$. The attempt frequencies for bulk and surface cavitation, $\kappa_{\mathrm{3D}}$ and $\kappa_{\mathrm{2D}}$, are obtained from fits of Eq. 1 to the simulation data; detail are given in the Supplementary Material, section S2.
III Results and Discussion
@@
III.1 Kinetic model for cavitation in water confined by surfaces with defects
@@
In the typical cavitation scenario, as studied in this work, water is confined by solid surfaces with a given contact angle that contain a finite number of surface defects. Upon application of a negative pressure, cavitation can proceed in bulk, at the defect-free surface parts or at the defects. Since we cannot presuppose the prevalence of one cavitation pathway over the others, we write the total cavitation rate as the sum of the bulk (3D), surface (2D) and defect (def) rates as
@@
where $N_{\mathrm{def}}$ is the number of identical surface defects. The model assumes that the different cavitation pathways do not interfere with each other. It will turn out that depending on the surface and defect properties, quantified by their respective contact angles, one cavitation pathway will always dominate strongly over the other pathways, such that we can construct effective kinetic phase diagrams that feature regions where bulk, surface or defect cavitation dominates.
III.2 Cavitation in bulk
CNT is a general framework for the description of cavitation Debenedetti (1996). For cavitation in bulk, one considers a spherical vapor bubble in a liquid with a free energy given by Caupin and Herbert (2006); Azouzi et al. (2013a); Menzl and Dellago (2016)
@@ III.2 Cavitation in bulk
where $\gamma$ is the liquid–vapor surface tension, $r$ the bubble radius and $p<0$ the negative pressure in the liquid. The first term represents the bubble surface free energy and the second term the work associated with the volume expansion. We have neglected the saturated vapor pressure, which is irrelevant compared to the large magnitude of cavitation pressures considered here. For simplicity, we have also neglected curvature effects on the surface tension, which only introduce minor corrections Menzl et al. (2016); Kanduč et al. (2020). The interplay between the two terms in Eq. 4 creates a free energy barrier for $p<0$ of
@@
$k_{\mathrm{3D}}=V\kappa_{\mathrm{3D}}\,e^{-\beta G^{*}_{\mathrm{3D}}}\,.$ (6)
The cavitation time follows from the rate as $\tau=k_{\mathrm{3D}}^{-1}$ Hänggi, Talkner, and Borkovec (1990) and is the mean time one has to wait for cavitation to occur. The exponential factor in Eq. 6 reflects the Arrhenius law Hänggi, Talkner, and Borkovec (1990). The pre-exponential factor splits into the volume $V$ and the attempt frequency density $\kappa_{\mathrm{3D}}$, which is an intensive property of the liquid that characterizes the barrier-less cavitation rate. It thus constitutes the cavitation speed limit, i.e., the maximum rate at which cavitation occurs ub te absence of a barrier Kubelka, Hofrichter, and Eaton (2004). The attempt frequency density, $\kappa_{\mathrm{3D}}$, accounts for viscosity and hydrodynamic memory effects in the surrounding liquid, as well as for corrections to CNT, such as the curvature dependence of the interfacial tension Acharya and Bagchi (2024). Given the exponential dependence of the cavitation time $\tau$ on $p$, it is preferred to consider the inverse relation. For a cubic volume $V=L^{3}$, the cavitation pressure follows from Eqs. 5 and 6 as  Caupin and Herbert (2006); Herbert, Balibar, and Caupin (2006)
@@
It is seen that $p_{\mathrm{3Dcav}}$ depends only logarithmically and thus rather weakly on system size $L$ and cavitation time $\tau$. Various methods for determining $\kappa_{\mathrm{3D}}$ have been proposed Blander and Katz (1975); Pettersen, Balibar, and Maris (1994). In a previous simulation study on bulk cavitation, we determined $\kappa_{\mathrm{3D}}=5\times 10^{11}\,\mathrm{ns^{-1}nm^{-3}}$ for the SPC/E water model Kanduč et al. (2020), this value will be employed in the current study. Using the same water model and the same methodology ensures consistency and enables a direct comparison between bulk, surface and defect cavitation rates. It is noteworthy that CNT (as given by Eq. 7) for SPC/E water predicts $p_{\mathrm{3Dcav}}$ values from --80 to --100 MPa for $L$ between 1 µm and 1 m and $\tau=1$ s. This value is slightly higher than previous predictions, primarily because the surface tension of SPC/E water ($\gamma_{\textrm{SPC/E}}\approx 55$ mN/m for a Lennard-Jones cutoff 0.9 nm Kanduč and Netz (2017)) is lower than the experimental value $\gamma_{\textrm{exp}}=72$ mN/m. Applying a correction factor of $(\gamma_{\textrm{exp}}/\gamma_{\textrm{SPC/E}})^{3/2}\approx 1.5$ on $p_{\mathrm{3Dcav}}$, as suggested by Eq. 7, yields good agreement with previous predictions Fisher (1948); Caupin (2005); Caupin and Herbert (2006); Caupin and Stroock (2013); Azouzi et al. (2013a). We here stick to the SCP/E water model because it reproduces most dynamic water properties quite well.
Experimental validation of the CNT bulk cavitation prediction for water is challenging, as the bulk cavitation pathway is preempted by diverse surface and defect cavitation pathways. Nevertheless, bulk cavitation occurs in liquids with considerably lower surface tensions, such as hydrocarbons. With surface tensions approximately three times smaller than water, the predicted cavitation pressures by CNT (Eq. 7) are reduced by a factor of $3^{3/2}\approx 5$, resulting in cavitation pressures of $-20$ to $-30$ MPa. This range agrees well with experimentally measured cavitation pressures in various organic liquids Galloway (1954); Apfel and Smith (1977); Ohde et al. (1993); Vinogradov and Pavlov (2000), providing validation of the CNT framework for bulk cavitation.
III.3 Cavitation at a smooth defect-free surface
@@
which is reduced compared to the bulk result $G_{\mathrm{3D}}^{*}$ in Eq. 5 by the geometric factor Debenedetti (1996); Caupin and Herbert (2006); Sear (2007)
@@
Figure 3: (A) Time-dependent pressure ($p$, top) and simulation box size ($L_{z}$, bottom) during a pressure ramp simulation with a pressure rate of $\dot{p}=-5\,\mathrm{MPa/ns}$ for a system with a surface contact angle of $\theta=97^{\circ}$. The presented trajectory corresponds to the final 200 ps prior to cavitation, identified as a 50 % increase of $L_{z}$ relative to its initial value $L_{z}^{0}=8\,\mathrm{nm}$. We find a cavitation pressure of $p_{\mathrm{cav}}^{*}=-93.9\,\mathrm{MPa}$. Vertical lines mark individual simulation segments at constant pressure. The inset shows consecutive snapshots of the cavitation event. (B) Cavitation configurations with $L_{z}/L_{z}^{0}=3/2$ for different contact angles. We also present trajectory visualizations for $\theta=45^{\circ}$ (upper left panel) (multimedia available online) and $\theta=135^{\circ}$ (lower right panel) (multimedia available online). (C) Simulation results for ${1/p_{\mathrm{cav}}^{*}}^{2}$ versus $\dot{p}$ for different contact angles. The lines are fits of Eq. 1 to the data, where $\kappa_{\textrm{2D}}$ is the only fitting parameter. (D) Results for $\kappa_{\textrm{2D}}$ (left scale) and the ratio $\xi=\kappa_{\mathrm{2D}}/\kappa_{\mathrm{3D}}$ (right scale) as a function of contact angle $\theta$. The solid black line is a fit of Eq. S13 in the Supplementary Material, section S3. For $\theta<52^{\circ}$, the fit does not serve as a reliable extrapolation and is denoted as a broken line.
@@
To determine $\kappa_{\textrm{2D}}$ from simulations, we use a self-assembled monolayer of alkyl molecules terminated by headgroups with adjustable electric dipole moments Kanduč, Schneck, and Netz (2014); Kanduč and Netz (2017). The dipole moment controls the wetting characteristics of the surfaces. We examine six different dipole moments, resulting in contact angles of $\theta=45^{\circ}$, $60^{\circ}$, $76^{\circ}$, $97^{\circ}$, $120^{\circ}$ and $135^{\circ}$, on surfaces measuring 9 nm $\times$ 10.4 nm in contact with water. Obtaining the cavitation time from simulations at constant negative pressure is impractical due to the excessive waiting time for cavitation to occur. Instead, we apply negative pressures that decrease linearly over time as $p(t)=\dot{p}t$ with different fixed pressure ramp rates $\dot{p}<0$, as used before for bulk cavitation  Kanduč et al. (2020) and described in the methods section.
@@
Fig. 3 summarizes our simulation results. Panel A illustrates a cavitation event on a hydrophobic surface with contact angle $\theta=97^{\circ}$, where a vapor bubble forms at the surface (see inset snapshots) due to the progressively decreasing pressure over time (upper graph). The simulation trajectory consists of a sequence of 20-ps-long simulations (delineated by gray vertical lines), each at a constant applied pressure that decreases in discrete steps of 0.1 MPa, which corresponds to a mean pressure rate of $\dot{p}=-5\,\mathrm{MPa/ns}$. The formation of the bubble is monitored by the box size normal to the surface $L_{z}$ (bottom graph). We define cavitation to take place when the box size $L_{z}$ increases by 50 % compared to its initial value $L_{z}^{0}$, which in this specific example yields a cavitation pressure $p_{\mathrm{2Dcav}}^{*}=-93.9\,\mathrm{MPa}$. It is important to note that the cavitation pressure in the pressure-ramp ensemble $p_{\mathrm{2Dcav}}^{*}$ is different from the cavitation pressure $p_{\mathrm{2Dcav}}$ in the constant-pressure ensemble given in Eq. 12, which normal pressure will be determined later.